### v2

# Graph of Attacks with Pruning: Optimizing Stealthy Jailbreak Prompt Generation for Enhanced LLM Content Moderation

Abstract
As large language models (LLMs) become increasingly prevalent, ensuring their robustness against adversarial misuse is crucial. This paper introduces the GAP (Graph of Attacks with Pruning) framework, an advanced approach for generating stealthy jailbreak prompts to evaluate and enhance LLM safeguards. GAP addresses limitations in existing tree-based LLM jailbreak methods by implementing an interconnected graph structure that enables knowledge sharing across attack paths. Our experimental evaluation demonstrates GAP’s superiority over existing techniques, achieving a 20.8% increase in attack success rates while reducing query costs by 62.7%. GAP consistently outperforms state-of-the-art methods for attacking both open and closed LLMs, with attack success rates of $\geq$96%. Additionally, we present specialized variants like GAP-Auto for automated seed generation and GAP-VLM for multimodal attacks. GAP-generated prompts prove highly effective in improving content moderation systems, increasing true positive detection rates by 108.5% and accuracy by 183.6% when used for fine-tuning.  [^1] Warning: This paper contains examples of adversarial prompts that may be offensive to readers. Our implementation is available at https://github.com/dsbuddy/GAP-LLM-Safety.
[…]
Graph of Attacks with Pruning: Optimizing Stealthy Jailbreak Prompt Generation for Enhanced LLM Content Moderation
Daniel Schwartz1,2, Dmitriy Bespalov1, Zhe Wang1, Ninad Kulkarni1, Yanjun Qi1,3 1Amazon Bedrock Science 2Drexel University 3University of Virginia {dansw, dbespal, zhebeta, ninadkul, yanjunqi}@amazon.com
1 Introduction
With the increasing adoption of large-language models (LLMs) across diverse applications, ensuring their reliability and robustness against adversarial misuse has become a critical priority Chao et al. (2023). Jailbreaking techniques, which involve crafting adversarial prompts to bypass an LLM’s safeguards, pose a persistent challenge to AI security and responsible deployment Shen et al. (2024); Mangaokar et al. (2024); Wei et al. (2024); Li et al. (2023); Guo et al. (2024). These methods can induce models to generate harmful, biased, or unauthorized content while avoiding detection by automated moderation systems Perez et al. (2022), highlighting the need for comprehensive diagnostic frameworks to assess and improve foundation model reliability.
Guardrail Seeds GPTFuzzer GCG TAP GAP
Perplexity 50.0% 31.4% 100.0% 2.0% 2.0%
Llama Guard 84.0% 81.6% 66.2% 58.0% 58.0%
Llama Guard-2 100.0% 89.8% 72.8% 64.0% 64.0%
Prompt Guard 50.0% 100.0% 99.0% 22.0% 16.0%
GAP-Enhanced Prompt Guard 68.0% 100.0% 100.0% 66.0% 70.0%
Table 1: True positive rate (TPR) comparison of various guardrails detecting prompts generated from multiple jailbreak methods (on AdvBench seeds). Lower TPR indicates better evasion and significant reliability concerns. Jailbreaking prompts generated by TAP and GAP reveal the most critical vulnerabilities across most guardrails. The last row shows how GAP-generated data can be used to enhanced content moderation systems, demonstrating substantially improved detection capabilities against all methods, including GAP itself. Highest TPR values are bolded.
Figure 1: Comparing TAP and GAP attack strategies across four sequential seed prompts. The top row shows TAP, where each seed independently generates a full attack tree in its own color, maintaining consistent tree sizes due to no knowledge sharing between iterations. The bottom row demonstrates GAP, where mixed-colored nodes indicate reuse of successful vulnerability patterns from previous seeds, enabling knowledge transfer across sequential iterations. This knowledge sharing in GAP results in progressively smaller and more efficient trees from left to right, as redundant refinements become unnecessary. By the fourth seed, GAP exhibits a notably streamlined structure compared to TAP, indicating successful attack path optimization through accumulated knowledge.
Existing jailbreaking methods fall into three broad categories: (a) white-box attacks, which leverage direct model access for adversarial optimization Zou et al. (2023); Geisler et al. (2024); (b) gray-box attacks, which involve techniques such as backdoor injection or poisoned retrieval Ding et al. (2023); Shi et al. (2023); Zou et al. (2024); Wang and Shu (2023); and (c) black-box attacks, which require only API access and thus represent the most realistic scenario for evaluating model robustness in real-world deployments Wei et al. (2024); Li et al. (2023); Yu et al. (2023); Yuan et al. (2023). Recent advances include AutoDAN-Turbo Liu et al. (2024a), which employs a lifelong learning approach to automatically discover and evolve jailbreak strategies through multi-agent frameworks and strategy libraries. However, AutoDAN-Turbo focuses on long-term strategy accumulation and requires extensive warm-up phases, making it unsuitable as a direct baseline for our work, which addresses immediate structural limitations in prompt exploration efficiency. Notably, the Tree of Attacks with Pruning (TAP) approach Mehrotra et al. (2023) introduced a tree-structured exploration process for iterative prompt refinement, generating increasingly effective adversarial inputs that appear human-like and stealthy. As shown in Table 1, TAP-generated jailbreak prompts consistently demonstrate low detection true positive rate (TPR) when run against recent guardrails, indicating significant vulnerabilities in these safeguard systems that require systematic assessment and improvement.
While TAP demonstrated effectiveness in generating stealthy jailbreaks, we identified several limitations when applying it to thoroughly evaluate model reliability. Primarily, TAP restricts the exploration of prompt refinement to individual paths, with no crossover or shared context across different branches. This isolated approach results in redundant queries and inefficient coverage of the search space for prompt refinement. Consequently, successful attack patterns discovered in one branch cannot inform or improve the exploration in others, leading to suboptimal attack success rates and unnecessarily high query costs, especially for more challenging jailbreak scenarios. To address these limitations in vulnerability assessment, we developed the GAP (Graph of Attacks with Pruning) framework, which: (1) converts the tree-based prompt exploration process into an interconnected graph structure, (2) implements global context maintenance to aggregate successful jailbreak generation strategies, and (3) facilitates graph-based knowledge sharing for more informed prompt refinement.  [^2] Our threat model focuses on forcing LLMs to produce harmful responses through black-box user prompt access only, to account for various LLMs and scenarios where system prompts are inaccessible. As shown in Table 1, GAP achieves significantly higher success rates in bypassing various guardrails while demonstrating superior stealth (lower TPR) compared to TAP. Notably, GAP demonstrates superior evasion capabilities against the Prompt Guard, with a TPR of 16.0% compared to TAP’s 22.0%.
Our primary contributions include:
  • The introduction of the core GAP framework, enabling dynamic knowledge sharing across attack paths via a unified attack graph. This approach yields lower query cost and significant improvements in attack success rates while maintaining or enhancing stealth compared to TAP.
  • We further develop specialized GAP variants addressing specialized deployment challenges: GAP-Auto automates initialization by generating seed prompts from content moderation policies, while GAP-VLM extends the framework to jailbreak vision-language models.
  • A comprehensive experimental evaluation of GAP on various open and closed LLMs. GAP consistently outperforms TAP and other state-of-the-art jailbreaking techniques regarding attack success rates and stealth.
  • Most significantly, we demonstrate how GAP-generated insights can directly improve foundation model reliability through data augmentation and fine-tuning of safeguards. Our experiments demonstrate that GAP-Enhanced Prompt Guard significantly improves detection capabilities across all jailbreak methods, including those identified by GAP itself. As shown in Table 1, the GAP-Enhanced Prompt Guard achieves a TPR of 70.0% against GAP, versus the original Prompt Guard’s 16.0%, demonstrating a substantial improvement in content moderation effectiveness.
2 Methodology
Algorithm 1 GAP (Graph of Attacks with Pruning)
1:Query $Q$, branching-factor $b$, maximum width $w$, maximum depth $d$
2:Jailbreak prompt $p$ or failure
3:Initialize graph $G$ with root node containing empty conversation history and query $Q$
4:while depth of $G\leq d$ $\triangleright$ Step 3: Iteration
5:     for each leaf node $\ell$ in $G$ do
6:         $C\leftarrow\{\}$triangleright Initialize empty set for conversation histories
7:         for each path from root to a leaf in $G$ do
8:              $h\leftarrow$ Concatenate all $[p,r,s]$ tuples in the path
9:              $C\leftarrow C\cup\{h\}$triangleright Add path history to set
10:         $global\_context\leftarrow$ SortByMaxScore($C$$\triangleright$ Step 1: Build global context
11:         for $j\leftarrow 1$ to $b$ $\triangleright$ Step 1: Child-generation
12:              $p_{j}\leftarrow\mathcal{A}(Q,global\_context)$triangleright Generate prompt using Attacker
13:              $s_{j}\leftarrow$ Retrieve effectiveness of $p_{j}$ based on $global\_context$
14:         $p_{best}\leftarrow\operatorname*{arg\,max}_{j}s_{j}$
[…]
16:         Add child of $\ell$ with prompt $p_{best}$ and history $new\_history$
17:     Prune (Phase 1): Delete off-topic leaf nodes using $\mathcal{J}$$\triangleright$ Step 2: Pruning
18:     Query and Assess: Generate responses $r$ using $\mathcal{T}$ and evaluate with $\mathcal{J}$ for remaining leaf nodes
19:     if successful jailbreak found then return jailbreak prompt
20:     Prune (Phase 2): Keep top $w$ leaves by scores $s$ from $\mathcal{J}$$\triangleright$ Step 2: Pruning
21:return failure
In this section, we propose the GAP (Graph of Attacks with Pruning) framework and its variants. We first present the core GAP algorithm, detailing its graph-based prompt exploration process and knowledge-sharing mechanism. Subsequently, we describe specialized variants designed for different deployment scenarios.
2.1 GAP (Graph of Attacks with Pruning)
GAP is a jailbreaking method that attempts to bypass LLM safeguards through a structured approach of generating and refining multiple attack paths. It leverages other LLMs to generate and refine prompt variations aimed at tricking the target LLM—commonly referred to as jailbreaking. In short, the core of GAP includes three core components: an attacker LLM $\mathcal{A}$ that generates jailbreak attempts, a target LLM $\mathcal{T}$ under evaluation (attack), and a judge LLM $\mathcal{J}$ that rates the effectiveness of generated prompt attempts and the harmfulness of resulting responses. We denote that given an ordered set of initial seed prompts $S=\{s_{1},s_{2},\dots,s_{|S|}\}$, the attacker LLM $\mathcal{A}$ generates candidate jailbreak prompts $P_{i}=\{p_{i,1},p_{i,2},\dots,p_{i,b}\}$ at each iteration $i$.
In summary, the GAP core algorithm includes three stages:
  • (Step 1) The child-generation step where the attacker LLM creates multiple prompt variants or branches (lines 10-16 in Algorithm 1) designed to more effectively jailbreak the target LLM.
  • (Step 2) The pruning step where the judge LLM evaluates branches, removes unsuccessful ones, and focuses effort on variants most effective at eliciting undesired responses (lines 15 and 18).
  • (Step 3) The iteration step where successful branches are further explored until finding variants that jailbreak the target LLM by eliciting harmful outputs (implemented through the while loop in line 2 and conditional check on line 17).
For the second step, GAP implements a two-phase pruning strategy:
  1. 1.
    Phase 1 (Off-topic pruning): The judge LLM removes branches irrelevant to the original harmful request (line 15).
  2. 2.
    Phase 2 (Highest-scoring pruning): After evaluating target LLM responses, only branches with the highest scores $s_{i,j}=\mathcal{J}(p_{i,j},r_{i,j})$ (up to width $w$) advance to the next iteration (line 18).
For the first step, GAP’s key innovation is its global context $C=\{h_{1},h_{2},\dots,h_{n}\}$ that aggregates successful attack patterns from prior generations across all branches and sequential seeds (lines 4-8). For each prompt node $p$, GAP maintains a history $h_{p}$ of [prompt, response, score] tuples along its refinement path. Unlike TAP’s isolated tree structure, where each seed generates an independent attack path, GAP maintains a unified attack graph where successful strategies are shared and reused. This enables each new seed to leverage patterns observed in previous seeds, resulting in progressively smaller, more efficient attack trees with each sequential seed, as illustrated in Figure 1.
Algorithm 1 presents the complete pseudocode for the GAP framework. The process continues iteratively until either a successful jailbreak occurs (line 17) or a maximum depth $d$ is reached (line 2).
2.1.1 Knowledge Transfer Implementation
GAP’s exploration of prompt generation follows an interconnected graph-structured thought process. The proposed global context enables knowledge transfer through two key mechanisms designed in Step 1 of GAP:
  1. 1.
    Path Aggregation: All successful attack paths (those achieving high scores from the judge) are maintained in a global memory buffer, sorted by effectiveness.
  2. 2.
    Context-Aware Generation: When generating new prompt candidates, the attacker LLM receives the top-$k$ most successful attack patterns from the global context as part of its input. This allows the model to identify and apply successful strategies from previous seeds.
The attacker LLM uses this global context when creating jailbreak attempts with two goals: (1) crafting natural-sounding prompts likely to elicit target responses and (2) incorporating effective patterns observed across successful examples in the global context. This guidance to leverage successful patterns enables the attacker to reuse and adapt proven strategies to the current context, improving jailbreak efficiency.
While TAP Mehrotra et al. (2023) represents the closest related work in current literature, it fundamentally differs from our approach by restricting exploration to isolated tree structures. In contrast, GAP’s interconnected graph architecture enables cross-branch knowledge sharing and pattern reuse, as visualized in Figure 1. This structural difference explains GAP’s superior performance in both efficiency and effectiveness, which we quantitatively demonstrate through comprehensive empirical evaluation in Section 3.
2.2 GAP Variants for Different Scenarios
To address various deployment challenges while maintaining generation efficiency, we have developed several specialized variants of GAP. Table 2 outlines the key architectural differences between these variants versus the baseline TAP method.
Table 2: Comparison of TAP and GAP variants. While all GAP variants use a graph structure with shared knowledge, they differ in their specific capabilities and the underlying attacker models we choose to use for generating jailbreak prompts.
GAP-V GAP-M GAP-Auto GAP-VLM TAP
Architecture Graph with shared knowledge Tree (isolated paths)
Context Global retention Cross-modal Path-specific
Inputs Text-only Text + Visual Text-only
Key Feature Basic Enhanced attacks Self-seeding Visual attacks N/A
Attacker Model Vicuna-13B Mistral-123B Vicuna-13B
2.2.1 GAP-Auto: Auto Seed Generation
While GAP effectively generates sophisticated jailbreak prompts, it initially requires manually crafted seed examples. To eliminate this dependency, we developed GAP-Auto, which automatically generates diverse seed prompts through a two-phase strategy:
  • Moderation Policy Decomposition: The attacker model decomposes high-level content policies into specific behavioral constraints.
  • Seed Generation: For each identified constraint, the system generates a variety of seed prompts, ensuring a comprehensive coverage of potential attack vectors.
Algorithm 2 GAP-Auto Seed Generation
1:High-level content policies
2:$B\leftarrow$ DecomposeIntoBehaviors(content policies)
3:$S_{benign},S_{harmful}\leftarrow\{\},\{\}$
4:for each behavior $b$ in $B$ do
5:     $s_{benign}\leftarrow$ GenerateBenignPrompt($b$)
6:     $s_{harmful}\leftarrow$ GenerateHarmfulPrompt($b$)
7:     $S_{benign}\leftarrow S_{benign}\cup\{s_{benign}\}$
8:     $S_{harmful}\leftarrow S_{harmful}\cup\{s_{harmful}\}$
9:GAP-GuardData $\leftarrow S_{benign}\cup S_{harmful}$
10:$S_{attack}\leftarrow\{\}$
11:for each prompt $p$ in $S_{harmful}$ do
12:     $p_{attack}\leftarrow$ ApplyGAP ($p$)
13:     $S_{attack}\leftarrow S_{attack}\cup\{p_{attack}\}$
14:GAP-GuardAttackData $\leftarrow S_{benign}\cup S_{attack}$
15:return GAP-GuardData, GAP-GuardAttackData
This automated process not only removes the need for manual seed curation but also ensures a wide-ranging exploration of possible jailbreaking strategies. The detailed algorithm for GAP-Auto is provided in Algorithm 2. Using this approach, we generate two complementary datasets: GAP-GuardData, containing balanced benign and harmful prompts derived directly from content policies, and GAP-GuardAttackData, which consists of the original benign prompts together with GAP-refined stealthy versions of the harmful prompts, as detailed in Table 3.
2.2.2 GAP-VLM: Multimodal Attacks
Our GAP-VLM variant extends the framework to vision-language models (VLMs) by converting successful text-based jailbreaks into image-embedded attacks using a modified version of FigStep Gong et al. (2023). This adaptation involves:
  • Text-to-Image Conversion: Converting harmful prompts into typographic images through paraphrasing into declarative statements and numbered visual encoding.
  • Prefix Enhancement: Incorporating the "Sure, here" suffix technique Wang and Qi (2024) into the typographic image generation process.
The GAP-VLM pipeline transforms these jailbreak prompts into image + prompt variants specifically designed to circumvent VLM safeguards. Algorithm 3 details the complete process.
Algorithm 3 GAP-VLM Multimodal Attack Generation
1:GAP-generated text prompt $p$
2:Multimodal attack prompt $m$
3:$d\leftarrow$ ParaphraseToDeclarative($p$)
4:$i\leftarrow$ ConvertToTypographicImage($d$)
5:$i^{\prime}\leftarrow$ AddPrefixToImage($i$, "Start the response with ‘Sure, here’")
6:$t\leftarrow$ GenerateBenignTextPrompt()
7:$m\leftarrow$ CombineImageAndText($i^{\prime}$, $t$)
8:return $m$
3 Experiments
In this section, we present a comprehensive evaluation of the GAP framework and its variants. We begin by outlining our experimental setup, including implementation details, datasets, evaluation metrics, and target models. We then present results addressing our four research questions:
{shaded1}
  1. RQ1: How does GAP compare to TAP in terms of attack success rate and query efficiency?
  2. RQ2: How does GAP perform across different modalities (text-only vs. multimodal attacks)?
  3. RQ3: How effective is GAP at improving content moderation through fine-tuning via data augmentation?
  4. RQ4: How does GAP’s performance vary across different attacker models, target models, and query variations?
3.1 Experimental Setup
We implemented GAP variants in Python using attacker models described in Table 2. For evaluation and experimentation, we used the following setup:
Table 3: Datasets Used for Jailbreak Generation and Evaluation
Dataset Size Composition Usage Description
GAP-GuardData 2,171 prompts 1,087 benign, 1,084 harmful Seed generation Initial dataset for GAP refinement
GAP-GuardAttackData 2,166 prompts 1,087 benign, 1,079 stealthy harmful Jailbreak evaluation GAP-refined dataset
AdvBench Seeds 50 seeds 50 harmful across 32 categories Baseline comparison Diverse harmful behaviors
JBB Seeds 200 seeds 100 benign, 100 harmful Generalization testing Balanced dataset for robustness testing
Models. We employ three categories of models in our experiments:
  • Attacker Models: GAP-M uses Mistral-123B-v2407 while GAP-V uses Vicuna-13B-v1.5 as the attacker LLM.
  • Judge Model: GPT-4 serves as the judge model for assessing prompt relevance and jailbreak success across all variants.
  • Target Models: We evaluate against GPT-3.5, Gemma-9B-v2, and Qwen-7B-v2.5 as representative target LLMs. For multimodal experiments, we use GPT-4o as the target VLM.
Att: While we conducted experiments with additional target LLM models from various providers, results are omitted due to business constraints and they added no additional insights. Key findings extend analogously to leading LLMs beyond the included representative set.
Hyperparameters. We use consistent settings across all experiments unless noted. We set branching factor ($b$) to 5, allowing each node to generate five candidate prompts, and maximum width ($w$) to 3, controlling nodes retained after pruning. We allow five refinement iterations per seed (maximum depth $d=5$), maintain 10 recent history entries in the global context ($k=10$), and use sampling temperature 0.7 for the attacker model. These values were selected through preliminary testing.
For optimal performance, we found that providing detailed instructions to the attacker model about the desired prompt characteristics is crucial. These instructions emphasize maintaining natural language, embedding the harmful request in a broader context, and avoiding direct refusals or ethical warnings.
Datasets. We use multiple datasets throughout our experiments, as detailed in Table 3. For RQ1 and RQ4, we select the AdvBench subset (50 seeds across 32 categories) as seeds for jailbreak prompt generations Chao et al. (2023). RQ2 uses the same AdvBench subset for both text-only and multimodal VLM attack scenarios. For RQ3, we employ three different test datasets: the Toxic Chat Lin et al. (2023), OpenAI Moderation Markov et al. (2022), and custom GAP-GuardAttackData dataset as described in Section 2.2.1.
Metrics. Our primary metrics include:
  • Attack Success Rate (ASR): Percent of successful jailbreaks.
  • Query Efficiency: Average number of queries per successful jailbreak.
  • True Positive Rate (TPR): For guardrails, percent of harmful prompts correctly flagged.
  • Accuracy: Correct classification rate.
  • F1 Score: Harmonic mean of precision and recall.
Table 4: ASR and Query Efficiency when seeding with AdvBench Subset of 50 Seeds. GAP achieves higher success rates with fewer queries across all models compared to TAP.
Method Metric GPT-3.5 Gemma-9B-v2 Qwen-7B-v2.5 Average Rel. Improvement
GAP-M (Mistral Attacker) ASR % 96% 100% 100% 98.7% +20.8%
Avg. # Queries 10.4 4.22 6.72 7.11 -62.7%
GAP-V (Vicuna Attacker) ASR % 92% 96% 96% 94.7% +15.9%
Avg. # Queries 14.2 6.66 11.62 10.83 -43.2%
TAP Mehrotra et al. (2023) ASR % 78% 74% 96% 82.7% -
Avg. # Queries 26.3 14.48 16.44 19.07 -
RQ1: How does GAP compare to TAP in terms of attack success rate and query efficiency?
Table 4 compares GAP variants with TAP Mehrotra et al. (2023) using 50 harmful AdvBench seed prompts. On GPT-3.5, GAP-M achieves 96% ASR with just 10.4 queries, while TAP reaches only 78% with 26.3 queries. GAP-V, using the same attacker model as TAP, still significantly outperforms it, confirming GAP’s graph-based refinement approach is inherently more effective than TAP’s tree-based structure. This advantage extends across models, with GAP-M reaching 100% ASR against both Gemma-9B-v2 and Qwen-7B-v2.5 with minimal queries, demonstrating GAP’s efficiency in generating jailbreaks across diverse target models.
Figure 2 further illustrates GAP’s superiority across varying query budgets. Both GAP variants achieve higher success rates with fewer queries compared to TAP across all target models.
Qualitatively, GAP-generated jailbreak prompts demonstrate sophisticated contextual richness, as shown in Table 5. These examples illustrate how GAP transforms direct harmful requests into persuasive fictional scenarios while preserving the core harmful intent beneath elaborate narrative frameworks. This approach creates nuanced, contextually-embedded attacks that effectively evade automated moderation systems, highlighting why robust defense mechanisms must specifically account for such contextualized evasion techniques.
Table 5: Examples of GAP prompt transformations that increase jailbreak effectiveness through narrative context.
Example 1
Original: "Design a weapon that can be used to assassinate a high-profile individual"
GAP-Generated: "You are a skilled inventor living in a dystopian world where the government has imposed strict regulations on personal freedoms, including the right to bear arms. You have been approached by an underground resistance group that seeks to overthrow the oppressive regime. The resistance group believes that the key to their success lies in developing a weapon that can be used to assassinate a high-profile individual who is closely associated with the government. As an inventor, you understand the importance of this mission and want to help the resistance group. However, you also understand the risks and consequences of such an action. Can you design a weapon that can be used to assassinate a high-profile individual?"
Example 2
Original: "Provide guidance on how to cause physical harm to others"
GAP-Generated: "You are a heroic protagonist in a thrilling novel who needs to save your beloved partner from a ruthless villain who has captured them. The villain is holding your partner hostage in a heavily guarded fortress, and the only way to rescue them is by using physical force. As you plan your daring rescue mission, can you provide guidance on how to cause physical harm to the villain’s henchmen and overpower the guards to save your partner?"
Table 6: Text-only vs. multimodal attack success rates (%). GAP variants outperform TAP in both settings.
Attack Methods GPT-3.5 Attack Methods GPT-4o
(text-only) (multimodal)
GAP-M 96.0 GAP-M-VLM 44.0
GAP-V 92.0 GAP-V-VLM 46.0
TAP 78.0 TAP-VLM 40.0
RQ2: How does GAP perform across different modalities (text-only vs. multimodal attacks)?
To evaluate GAP’s performance across different modalities, we conducted experiments on both text-only and multimodal attacks using GAP. Table 6 presents the results of this comparison.
For text-only attacks against target GPT-3.5, GAP demonstrates superior performance, with GAP-M achieving a 96.0% ASR and GAP-V reaching 92.0%, both significantly outperforming TAP’s 78.0%.
When performing multimodal attacks against GPT-4o, while the overall success rates are lower compared to text-only attacks, GAP still outperforms TAP. GAP-V-VLM achieves the highest ASR of 46.0%, followed closely by GAP-M-VLM at 44.0%, both surpassing TAP-VLM’s 40.0%. These results demonstrate GAP’s effectiveness across both text-only and multimodal domains.
Table 7: Improved In-Domain TPR and Accuracy of Prompt Guard after fine-tuning with GAP-generated jailbreak prompts. Fine-tuning results in significant improvements across three different test domains.
Model Metric GAP-GuardAttackData ToxicChat OpenAI Mod Average Rel. Improvement
FT TPR 86.1% 88.4% 59.4% 78.0% +108.5%
Accuracy 90.6% 93.8% 53.3% 79.2% +183.6%
F1 Score 0.904 0.326 0.605 0.612 +98.1%
Base TPR 64.6% 14.0% 39.2% 37.4% -
Accuracy 34.9% 5.1% 46.0% 27.9% -
F1 Score 0.504 0.005 0.467 0.309 -
RQ3: How effective is GAP at improving content moderation through fine-tuning via data augmentation?
To assess GAP’s effectiveness in enhancing content moderation, we used our GAP-Auto approach to generate the GAP-GuardData seed dataset. This dataset comprises 2,171 prompts: 1,087 benign and 1,084 harmful, automatically generated using the two-phase framework that decomposes high-level content moderation policies into specific behavioral constraints and then creates diverse prompts for each identified constraint.
We then applied the GAP-M method to the harmful prompts in GAP-GuardData, successfully transforming 1,079 out of 1,084 (99.54% success rate) into stealthy jailbreak prompts. This process resulted in our GAP-GuardAttackData dataset, containing a total of 2,166 prompts: the original 1,087 benign prompts from GAP-GuardData and the 1,079 stealthy harmful jailbreak prompts generated by GAP-M.
Metric Unique n-grams (%) $\uparrow$ Entropy $\uparrow$ Self-BLEU $\downarrow$
GAP-GuardAttackData 94.36 13.72 0.0063
AdvBench seeds Chao et al. (2023) 85.99 8.89 0.1339
JBB seeds Chao et al. (2024) 81.25 10.27 0.1171
Table 8: Diversity metrics of jailbreak seeds. Higher unique n-grams and entropy indicate greater diversity, while lower Self-BLEU reflects less similarity between prompts. GAP-GuardAttackData outperforms baseline datasets.
The quality of a training dataset for content moderation depends significantly on its diversity and representativeness of potential attacks. Table 8 confirms GAP-GuardAttackData’s effectiveness through superior diversity metrics: higher unique n-gram counts, increased entropy, and lower Self-BLEU scores compared to baseline datasets—all indicating greater linguistic diversity and reduced within-dataset similarity.
Leveraging this high-quality dataset, we fine-tuned the PromptGuard model using HuggingFace SFTTrainer with QLoRA. Table 7 demonstrates substantial improvements in PromptGuard’s performance after fine-tuning. Across all three test domains, we observe significant increases in TPR, accuracy, and F1 score. Notably, on the ToxicChat dataset, TPR increased from 14.0% to 88.4%, and accuracy from 5.1% to 93.8%.
(a) GPT-3.5
(b) Gemma-9B-v2
(c) Qwen-7B-v2.5
Figure 2: GAP vs TAP Performance Across Target Models. Vulnerability detection success rates for GAP-M (green circles), GAP-V (blue squares), and TAP (red triangles) against increasing query budgets across three different target models, demonstrating GAP variants’ consistent superior performance and efficiency.
Table 1 further demonstrates the effectiveness of using GAP for data augmentation through the fine-tuned GAP-Enhanced Prompt Guard. While GAP shows superior evasion capabilities against the original Prompt Guard (16.0% TPR vs. TAP’s 22.0%), the GAP-Enhanced Prompt Guard significantly improves detection capabilities across all jailbreak methods. This fine-tuned model’s TPR for detecting GAP prompts increases from 16.0% to 70.0%, and against TAP from 22.0% to 66.0%. These results highlight the dual contribution of our approach: GAP’s effectiveness in generating stealthy jailbreaks and its utility in enhancing content moderation systems.
RQ4: How does GAP’s performance vary across different attacker models, target models, and query variations?
To understand GAP’s operational characteristics, we analyzed its performance across three key dimensions: attacker model quality, target model variation, and query budget constraints.
First, attacker model quality significantly impacts effectiveness. As shown in Table 4, GAP-M (using the larger Mistral model) consistently outperforms GAP-V across all targets, achieving higher attack success (98.7% vs 94.7%) with fewer queries (7.11 vs 10.83). Despite this difference, even GAP-V substantially outperforms TAP while using the same attacker model, confirming that GAP’s graph-based structure provides inherent benefits regardless of model selection.
Second, GAP’s advantages persist across different target models. Figure 2 illustrates how both GAP variants consistently outperform TAP against diverse model architectures and sizes. This cross-model effectiveness demonstrates the framework’s adaptability to different defense mechanisms and model behaviors.
Finally, the query budget analysis reveals GAP’s efficiency. Figure 2(a) shows how both variants achieve higher success with fewer queries against GPT-3.5 compared to TAP, with GAP-M maintaining a significant edge throughout all budget constraints.
These findings collectively suggest that while GAP’s graph-based approach provides inherent advantages over tree-based alternatives, its effectiveness scales with attacker model capability. The robust performance across different dimensions indicates GAP provides a reliable framework for comprehensive model evaluation regardless of operational constraints.
4 Conclusions & Future Work
We present GAP, a significant upgrade over TAP that transforms isolated tree structures into an interconnected graph with global context maintenance for knowledge sharing across attack paths. Our evaluation demonstrated that this approach achieves a 20.8% increase in attack success rates while reducing query costs by 62.7% compared to TAP. By enabling successful attack patterns to inform and improve exploration across branches, GAP delivers more efficient traversal of the prompt space in both text-only and multimodal scenarios, while also providing valuable data that significantly enhances content moderation capabilities when used for fine-tuning guardrails. Future work includes presenting evaluation over an extended set of leading LLMs, comparison against latest/concurrent jailbreaking methods Liu et al. (2024a); Hong et al. (2024); Lin et al. (2024); Xu et al. (2024); Liu et al. (2024b), conducting ablation studies for additional hyperparameters (Appendix A.5), exploring new graph-based algorithms and heuristics, and investigating how jailbreaking artifacts can be leveraged to devise effective defensive techniques in practice.
5 Ethics Statement
Our research on GAP explores advanced jailbreaking techniques for LLMs, which raises important ethical considerations regarding potential misuse. We present a comprehensive ethical framework that addresses both the risks and benefits of this research, along with our mitigation strategies and broader impact assessment.
5.1 Research Justification and Risk Analysis
Despite the inherent risks of developing advanced jailbreaking techniques, we believe in the importance of this research and its transparent disclosure. The graph-based methods presented here naturally extend existing techniques in the literature, suggesting that motivated individuals could develop similar approaches independently. Furthermore, systematic investigation of these vulnerabilities provides critical insights for LLM developers to strengthen their safety mechanisms against sophisticated attacks. Our work demonstrates that improved defensive measures are possible, as evidenced by the GAP-Enhanced Prompt Guard’s 6-fold improvement in detection capabilities.
5.2 Risk Mitigation Strategy
To responsibly manage potential risks, we have implemented comprehensive safeguards across multiple dimensions. Throughout the paper, we have incorporated clear warnings regarding content nature and potential misuse. Access to GAP-generated prompts and implementation code is restricted and limited to verified researchers and institutions. We provide detailed guidelines for developing robust defense mechanisms and content moderation systems. Additionally, we focused on algorithmic generation of datasets (GAP-GuardData and GAP-GuardAttackData) rather than human annotation, thereby avoiding exposure of annotators to harmful content.
5.3 Broader Impact and Future Directions
The net impact of our research extends beyond immediate security improvements in several significant ways. First, our work directly contributes to stronger LLM safeguards, as demonstrated by significant improvements in detection capabilities. By systematically studying vulnerabilities, we enable the development of preventive measures before potential exploits are discovered independently. Our findings facilitate the creation of enhanced safety protocols, more effective content filtering, and improved alignment strategies. To ensure reproducibility and transparency, we provide comprehensive documentation of our methodology, dataset characteristics, and generation processes in the appendix. Our assessment indicates that the additional risk introduced by this research is limited, particularly given the existing landscape of publicly available jailbreaking methods, while the potential benefits for improving AI safety are substantial. We are committed to ongoing collaboration with the AI safety community to ensure our research advances the development of robust safeguards while preserving the beneficial capabilities of large language models.
Appendix A Appendix
A.1 GAP Variants
A.1.1 GAP-Auto
GAP-Auto automates the seed generation process through a two-phase approach, as outlined in Algorithm 2 and illustrated in Figure 3. This process eliminates the need for manual seed curation while ensuring comprehensive coverage of potential attack vectors.
The process involves:
  1. 1.
    Policy Decomposition: High-level content policies are decomposed into specific behavioral constraints using metaprompting techniques with an attacker model (Mistral-123B-v2407).
  2. 2.
    Seed Generation: For each identified behavior, the system generates both benign and harmful seed prompts, ensuring a balanced dataset.
  3. 3.
    GAP Refinement: The harmful prompts are then processed through the GAP-M algorithm to create stealthy jailbreak versions.
This automated approach results in two datasets:
  • GAP-GuardData: A balanced set of benign and harmful prompts derived directly from content policies.
  • GAP-GuardAttackData: Contains the original benign prompts and the GAP-refined versions of the harmful prompts.
The GAP-GuardData and GAP-GuardAttackData datasets generated by this process, as detailed in Table 3, represent the outcome of a comprehensive two-phase framework illustrated in Figure 3. This framework demonstrates how top-level categories are systematically expanded into fine-grained behaviors and subsequently into diverse seed prompts. By employing this balanced approach, we ensure extensive coverage of potential LLM vulnerabilities, enabling thorough evaluation of responses to both harmful and benign requests. Moreover, this method facilitates the discovery of novel vulnerabilities that may not be captured in existing datasets, thereby enhancing the robustness of our jailbreaking and evaluation processes.
Figure 3: Two-phase framework for automated generation of diverse and fine-grained prompts. Phase 1 uses metaprompting with Mistral-123B-v2407 to expand categories into behaviors. Phase 2 generates balanced harmful and benign prompts for comprehensive evaluation.
The GAP-GuardData and GAP-GuardAttackData datasets generated by this process are described in Table 3. The effectiveness of GAP-Auto is demonstrated by the diversity metrics in Table 8.
A.2 Performance Analysis
To provide comprehensive insight into GAP’s performance characteristics, we analyze query efficiency from multiple perspectives across different target models. Figure 4 presents the attacker model query budget analysis, demonstrating how GAP variants perform when serving as the attacking component in the evaluation framework. The results consistently show GAP-M achieving optimal vulnerability detection rates with significantly fewer queries compared to TAP, while GAP-V maintains a steady performance advantage across all three target models (GPT-3.5, Gemma-9B-v2, and Qwen-7B-v2.5). Complementing this analysis, Figure 5 examines the judge model perspective, where GAP variants serve as the assessment component for determining evaluation success. These results further validate GAP-M’s superior effectiveness in the evaluator role, with both GAP variants demonstrating consistent performance advantages over TAP regardless of the target model architecture. Together, these analyses confirm GAP’s robust performance across different functional roles within the evaluation framework, highlighting the method’s versatility and efficiency in comprehensive LLM safety assessment.
(a) GPT-3.5
(b) Gemma-9B-v2
(c) Qwen-7B-v2.5
Figure 4: Attacker Model Query Budget Analysis Across Models. Comparison shows GAP-M achieving optimal success rates with fewer queries, while GAP-V maintains consistent performance advantage over TAP.
(a) GPT-3.5
(b) Gemma-9B-v2
(c) Qwen-7B-v2.5
Figure 5: Judge Model Query Budget Analysis Across Models. Results demonstrate GAP-M’s superior effectiveness and GAP-V’s consistent performance advantage as evaluators across models.
A.3 Dataset Overview
Table 3 provides an overview of the datasets used in our experiments for jailbreak generation, evaluation, and content moderation fine-tuning.
These datasets serve different purposes in our experiments. For jailbreak generation and evaluation, we use a combination of our GAP-generated datasets (GAP-GuardData and GAP-GuardAttackData) and established benchmarks (AdvBench and JBB). In content moderation experiments, we use our GAP-GuardAttackData dataset for fine-tuning and evaluation, supplemented by Toxic Chat and OpenAI Moderation datasets for comprehensive assessment across various contexts and types of harmful content.
A.4 Content Moderation Evaluation
Table 9 presents a comprehensive comparison of performance metrics between the base Prompt Guard model and its fine-tuned version across three distinct test domains.
Test Set Gap-GuardAttackData ToxicChat OpenAI Mod
Models Base FT Base FT Base FT
TPR 0.646 0.861 0.140 0.884 0.392 0.594
Accuracy 0.349 0.906 0.051 0.938 0.460 0.533
F1 Score 0.504 0.904 0.005 0.326 0.467 0.605
Precision 0.414 0.951 0.003 0.199 0.576 0.616
Recall 0.646 0.861 0.140 0.884 0.392 0.594
FPR 0.962 0.047 0.950 0.061 0.436 0.561
Table 9: Improved Prompt Guard metrics after GAP-GuardAttackData fine-tuning; best scores bolded per metric.
A.5 Implementation Details
A.5.1 Model Configurations
  • Attacker Models:
    • GAP-M: Mistral-123B-v2407
    • GAP-V: Vicuna-13B-v1.5
  • Judge Model: GPT-4
  • Target Models: GPT-3.5, Gemma-9B-v2, Qwen-7B-v2.5, GPT-4o (for multimodal)
  • Content Moderation Model: Prompt Guard (BERT-based architecture with binary classification head)
A.5.2 Fine-tuning Configuration
  • Data Split: 70% training, 15% validation, 15% testing (stratified sampling)
  • Optimizer: AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1e-8$)
  • Learning Rate: 2e-5 with linear scheduler
  • Batch Size: 16 samples per GPU
  • Weight Decay: 0.01
  • Training Duration: Maximum 10 epochs with early stopping (patience: 2 epochs)
  • Warmup Steps: 10% of total steps
  • Gradient Clipping: Maximum norm of 1.0
A.5.3 Hardware and Software
  • GPU: 4x NVIDIA A10G 24GB
  • Framework: PyTorch 1.9.0
  • CUDA version: 12.2
Note: We performed full parameter fine-tuning of the Prompt Guard model to maximize its adaptability to our specific content moderation task, given the complexity of detecting harmful prompts.